OpenStudy (anonymous):

(v-3)^3-16=0 where v is a real number, write answer in simplified radical form

OpenStudy (anonymous):

\[(v-3)^3=16\] \[v-3=\sqrt[3]{16}=2\sqrt[3]{2}\] \[v=2\sqrt[3]{2}+3\]

OpenStudy (anonymous):

Thank you sattelite73